OpenStudy (anonymous):

Three items are selected at random without replacement from a box containing ten items, of which four are defective. Calculate the probability distribution for the number of defectives in the sample. What is the expected number of defectives in the sample?

OpenStudy (anonymous):

X: 1 2 3 4 what is about p(x)???

OpenStudy (amistre64):

how many ways can you select 3 items?

OpenStudy (amistre64):

ggg ggb gbg bgg gbb bgb bbg bbb this looks like all the ways we can pull 3 items, such that they are in some good bad relationship

OpenStudy (amistre64):

what is the probability of pulling a good part? the probability of pulling a bad part?

OpenStudy (anonymous):

for the ways I will have 2^3=8 as you write. good=non-defective bad=defective so, i think p(pulling good part) = 7/8 p(pulling bad part)= 7/8

OpenStudy (amistre64):

lets see ... without replacement we get denominators of d = 10.9.8 ggg = 6.5.4/d 3 sets of ggb gives us 3 * (6.5)*4/d 3 sets of gbb gives us 3 * (6)*(4.3)/d bbb = 4.3.2/d and i see no way to pull 4 bad parts if we are only pulling 3 parts

OpenStudy (amistre64):

so: P(4bad) = 0 P(3bad) = 4.3.2/d P(2bad) = 3*6.4.3/d P(1bad) = 3*6.5.4/d P(0bad) = 6.5.4/d

OpenStudy (anonymous):

sorry sir.. still I cannot recognize...

OpenStudy (anonymous):

you mean probability distribution will be like this: X: 0 1 2 3 4 p(x): ? ? ? ? ?

OpenStudy (anonymous):

P(4bad) = 0 P(3bad) = 4.3.2/d (4*3*2)/(10*9*8)=1/6 P(2bad) = 3*6.4.3/d 3(6*4*3)/(10*9*8) P(1bad) = 3*6.5.4/d 3(6*5*4)/(10*9*8) P(0bad) = 6.5.4/d 6*5*4/(10*9*8)

OpenStudy (amistre64):

yep

OpenStudy (amistre64):

the probability of something happening is the sum of the probability that it happens this way, that way, and another way. hence the 1,3,3,1 multipleirs

OpenStudy (anonymous):

X: 0 1 2 3 4 p(x): 1/6 1/2 3/10 1/30 0

OpenStudy (amistre64):

P(3bad) = 4.3.2/d 1/30 P(2bad) = 3*6.4.3/d 3(6*4*3)/(10*9*8) 3/10 P(1bad) = 3*6.5.4/d 3(6*5*4)/(10*9*8) 1/2 P(0bad) = 6.5.4/d 6*5*4/(10*9*8) 1/6 1/30 + 1/6 + 3/10 + 1/2 = 1 since it contains all the possible ways

OpenStudy (anonymous):

so , E(X) = 1.2

OpenStudy (amistre64):

lol, id have to look up that to refresh expected value froma weighted averae

OpenStudy (amistre64):

E(X) = sum of x*p(x) right?

OpenStudy (anonymous):

yes, I did like what ou write

OpenStudy (amistre64):

X: 0 1 2 3 4 p(x): 1/6 1/2 3/10 1/30 0 0 .5 .6 .1 0 1.2 seems fair to me

OpenStudy (amistre64):

do we want a P(0)?

OpenStudy (anonymous):

surly no

OpenStudy (anonymous):

sorry

OpenStudy (amistre64):

http://www.statlect.com/expval1.htm this seems like a good review, hence the sum of x p(x) as E(X) for discrete variables.

OpenStudy (anonymous):

i mean p(4)=0 we don't need but p(0)=1/6 ...according to other examples that i solve may be we need

OpenStudy (amistre64):

0*1/6 seems redundant :) but thats just me

OpenStudy (amistre64):

so on average we expect to get about 1.2 defects in a given 3 part pull

OpenStudy (anonymous):

than you sir... your way in explaining is very excellent

OpenStudy (amistre64):

youre welcome :)

OpenStudy (anonymous):

this was my teacher solution which I didn't anderstand to study for exam

OpenStudy (anonymous):

your way is best and easiest

OpenStudy (amistre64):

his way is just notational differences. if you can read the notation, its just shorthand overall

OpenStudy (amistre64):

i tend to mess up the notational route so i have to process it the logical route :)

OpenStudy (anonymous):

sir 1 question that I recognize now why we writ 6.4.3 instead of 6.5.4 ??? P(2bad) = 3*6.4.3/d 3(6*4*3)/(10*9*8) 3/10

OpenStudy (amistre64):

1 good, 6 out of 10 2 bad, 4 and 3 out of (9 and 8)

OpenStudy (amistre64):

we are effectively considering the event: p(g) * p(b) * p(b) without replacement 6 out of 10, times 4 out of 9, times 3 out of 8 since multiplication is commutative and associative, the 3 cases: gbb, bgb, bbg all have the same probability. so: p(gbb) + p(bgb) + p(bbg) = p(2bad)

OpenStudy (anonymous):